OpenStudy (anonymous):

Help me please! (3/a^3) + (9a/a^2) + (2/a^3)

OpenStudy (anonymous):

\[\frac{ 3 }{ a^3 } +\frac{ 9a }{ a^2 } + \frac{ -2 }{a^3 }\]

OpenStudy (anonymous):